OpenStudy (anonymous):

find the mass of a ball given x^2 + y^2 + z^2 = a^2 if density at any point is proportional to distance from z axis

OpenStudy (anonymous):

use \[\int\limits \int\limits \int\limits k \sqrt{x ^{2}+y ^{2}}dV\]